How do I divide a number by 2sin(theta)?
As an example from my book: 300 N/2 sin (5) = 1700 N
When I do it on the calc: 300/(sin(5))^2, I get 326. If I do, 300/2sin(5) then I get -143.
I have no clue what to do! It has been many, many years since I had algebra and I have never had a trig class. I am currently in physics and quickly getting behind due to my shaky math. Any help would be GREATLY appreciated. Thank you!

Answer by Alan3354(69443)

How do I divide a number by 2sin(theta)?
As an example from my book: 300 N/2 sin (5) = 1700 N
When I do it on the calc: 300/(sin(5))^2, I get 326. If I do, 300/2sin(5) then I get -143.
------------------
300/(2sin(5)) = 150/sin(5)
=~ 1721 (if it's 5 degs, not 5 radians)
---
That's close to 1700, but I don't know why it says 1700, not 1721.
I see no reason to square the sine.